# OneIG-Bench: Omni-dimensional Nuanced Evaluation for Image Generation

### v2

@@ 4.2 Results and Analysis
We evaluate a range of well-known image generation models on our benchmark, including unified multimodal models (Janus-Pro [13], BLIP3-o [8], BAGEL [18]), open-source models (Stable Diffusion 1.5 [55], Stable Diffusion XL [46], Stable Diffusion 3.5 [60], Flux.1-dev [38], CogView4 [76], SANA [72], Lumina-Image 2.0 [47], and HiDream-I1-Full [29]) with A800 GPUs, as well as closed-source models (Imagen3 [36], Recraft V3 [66], Kolors 2.0 [64], Seedream 3.0 [24], Imagen4 [17] and GPT-4o [45]). To present a comprehensive comparison, we aggregate evaluation metrics across multiple dimensions, as summarized in Table 2. We define the sets of images generated based on the OneIG-Bench prompt categories General Object $\mathcal{O}$, Portrait $\mathcal{P}$, Anime and Stylization $\mathcal{A}$ (prompts without stylization), $\mathcal{S}$ (prompts with stylization), Text Rendering $\mathcal{T}$, Knowledge and Reasoning $\mathcal{KR}$ and Multilingualism $\mathcal{L}$. A more detailed, fine-grained analysis of individual dimensions on OneIG-Bench follows in the subsequent sections and the appendix.
@@ 4.2.2 Text Rendering
Notably, Recraft V3 [66] consistently achieves excellent edit distance performance across all prompt lengths, with its values for long prompts being less than half of the second-best result. However, its performance in completion ratio and word accuracy count is relatively less impressive. This discrepancy may be attributed to its layout-first strategy [65], in which a layout is generated prior to text insertion. This approach effectively reduces severe errors and prevents chaotic outputs by decomposing the original text rendering task into several relatively simpler subtasks, thereby significantly enhancing the reliability of text rendering in the final images.
@@ A.5 Visualization Results
We first give the best normalized polar visualization of SOTA methods on OneIG-Bench in Table 12. It can be observed that the closed-source methods outperform the other two categories of methods overall. And we selected the top five methods based on their overall performance across non-style metrics and showcased representative examples for each. For the style dimension, we further selected images from three finer-grained subcategories to illustrate the results. In all visualizations, the methods are ordered from left to right according to their performance ranking, from highest to lowest. In the following figures, each image tile is labeled in the top-left corner with the score achieved by the corresponding method under the current evaluation metric.
@@
Table 12: Best normalized polar visualization of SOTA methods on OneIG-Bench. Anti-Clockwisely, Alignment: NP, T&P, Short, Medium, Long; Text: ED-Short, ED-Medium, ED-Long, CR-Short, CR-Medium, CR-Long, WAC-Short, WAC-Medium, WAC-Long; Reasoning: Geography,   Computer Science,   Biology,   Mathematics,   Physics,   Chemistry,   Common Sense; Style: Traditional, Media, Anime; Diversity: NP, T&P, Short, Medium, Long. The absolute average values of each evaluation metric are list on legends.
@@
BAGEL+CoT Stable Diffusion 1.5 Stable Diffusion XL
@@
Table 13: [Continued] Best normalized polar visualization of SOTA methods on OneIG-Bench. Anti-Clockwisely, Alignment: NP, T&P, Short, Medium, Long; Text: ED-Short, ED-Medium, ED-Long, CR-Short, CR-Medium, CR-Long, WAC-Short, WAC-Medium, WAC-Long; Reasoning: Geography,   Computer Science,   Biology,   Mathematics,   Physics,   Chemistry,   Common Sense; Style: Traditional, Media, Anime; Diversity: NP, T&P, Short, Medium, Long. The absolute average values of each evaluation metric are list on legends.
@@
Figures 1112, and 13 show that GPT-4o [45] performs well across most styles, though it struggles with specific ones especially some anime styles. Stable Diffusion 1.5 [55], despite its lower visual quality, effectively captures traditional style features. Although Imagen4 [17] does not achieve a high overall score in style, it performs notably well in media and anime styles. It is worth noting that unified multimodal methods, such as BAGEL+CoT [18], BAGEL [18], and Janus-Pro [13], exhibit particularly strong performance in the anime style. Overall, Seedream 3.0[24] and SANA-1.5 4.8B (PAG)[72] also demonstrate strong stylistic consistency, ranking just behind GPT-4o.

### v3

@@ 4.2 Results and Analysis
We evaluate a range of well-known image generation models on our benchmark, including unified multimodal models (Janus-Pro [13], BLIP3-o [8], BAGEL [18], Show-o2 [75], OmniGen2 [71]), open-source models (Stable Diffusion 1.5 [55], Stable Diffusion XL [46], Stable Diffusion 3.5 [60], Flux.1-dev [38], CogView4 [78], SANA [73], Lumina-Image 2.0 [47], and HiDream-I1-Full [29]) with A800 GPUs, as well as closed-source models (Imagen3 [36], Recraft V3 [66], Kolors 2.0 [64], Seedream 3.0 [24], Imagen4 [17] and GPT-4o [45]). To present a comprehensive comparison, we aggregate evaluation metrics across multiple dimensions, as summarized in Table 2. We define the sets of images generated based on the OneIG-Bench prompt categories General Object $\mathcal{O}$, Portrait $\mathcal{P}$, Anime and Stylization $\mathcal{A}$ (prompts without stylization), $\mathcal{S}$ (prompts with stylization), Text Rendering $\mathcal{T}$, Knowledge and Reasoning $\mathcal{KR}$ and Multilingualism $\mathcal{L}$. A more detailed, fine-grained analysis of individual dimensions on OneIG-Bench follows in the subsequent sections and the appendix.
@@ 4.2.2 Text Rendering
Notably, Recraft V3 [66] consistently achieves excellent edit distance performance across all prompt lengths, with its values for long prompts showing a large gap from the second-best result. However, its performance in completion ratio and word accuracy count is relatively less impressive. This discrepancy may be attributed to its layout-first strategy [65], in which a layout is generated prior to text insertion. This approach effectively reduces severe errors and prevents chaotic outputs by decomposing the original text rendering task into several relatively simpler subtasks, thereby significantly enhancing the reliability of text rendering in the final images.
@@ A.2 Implementation
Show-o2-1.5B [75] 1.5B 5.0 432 $\times$ 432 50
Show-o2-7B [75] 7B 5.0 432 $\times$ 432 50
OmniGen2 [71] 3B+4B 4.0 1024 $\times$ 1024 50
@@ A.5 Visualization Results
We first give the best normalized polar visualization of SOTA methods on OneIG-Bench in Table 12. It can be observed that the closed-source methods outperform the other two categories of methods overall. And we selected some methods based on their overall performance across non-style metrics and showcased representative examples for each. For the style dimension, we further selected images from three finer-grained subcategories to illustrate the results. In all visualizations, the methods are broadly selected based on overall performance. In the following figures, each image tile is labeled in the top-left corner with the score achieved by the corresponding method under the current evaluation metric.
@@
Table 12: Best normalized polar visualization of SOTA methods on OneIG-Bench. Anti-Clockwisely, Alignment: NP, T&P, Short, Middle, Long; Text: ED-Short, ED-Middle, ED-Long, CR-Short, CR-Middle, CR-Long, WAC-Short, WAC-Middle, WAC-Long; Reasoning: Geography,   Computer Science,   Biology,   Mathematics,   Physics,   Chemistry,   Common Sense; Style: Traditional, Media, Anime; Diversity: NP, T&P, Short, Middle, Long. The absolute average values of each evaluation metric are list on legends.
@@
BAGEL+CoT Show-o2-1.5B(432) Show-o2-7B(432)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
OmniGen2 Stable Diffusion 1.5 Stable Diffusion XL
@@
Table 13: [Continued] Best normalized polar visualization of SOTA methods on OneIG-Bench. Anti-Clockwisely, Alignment: NP, T&P, Short, Middle, Long; Text: ED-Short, ED-Middle, ED-Long, CR-Short, CR-Middle, CR-Long, WAC-Short, WAC-Middle, WAC-Long; Reasoning: Geography,   Computer Science,   Biology,   Mathematics,   Physics,   Chemistry,   Common Sense; Style: Traditional, Media, Anime; Diversity: NP, T&P, Short, Middle, Long. The absolute average values of each evaluation metric are list on legends.
@@
Figures 1112, and 13 show that GPT-4o [45] performs well across most styles, though it struggles with specific ones especially some anime styles. Stable Diffusion 1.5 [55], despite its lower visual quality, effectively captures traditional style features. Although Imagen4 [17] does not achieve a high overall score in style, it performs notably well in media and anime styles. It is worth noting that unified multimodal methods, such as BAGEL+CoT [18], OmniGen2 [71], BAGEL [18] and Janus-Pro [13], exhibit particularly strong performance in the anime style. Overall, Seedream 3.0[24] and SANA-1.5 4.8B (PAG)[73] also demonstrate strong stylistic consistency, ranking just behind GPT-4o.